The CEO of a battery manufacturing company claimed that their batteries would last an average of 280 hours under normal use. A researcher randomly selected 20 batteries from the production line and tested them. The tested batteries had a mean life span of 250 hours with a standard deviation of 40 hours. Do we have enough evidence to suggest that the claim of an average of 280 hours is false?


1
Expert's answer
2022-04-27T17:05:51-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=280"

"H_1:\mu\not=280"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha = 0.05," "df=n-1=20-1=19" degrees of freedom, and the critical value for a two-tailed test is "t_c = 2.093024"

The rejection region for this two-tailed test is"R = \{t: |t| > 2.023024\}."

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{250-280}{40/\sqrt{20}}\approx-3.354102"

Since it is observed that "|t| = 3.354102 >2.093024= t_c," it is then concluded that the null hypothesis is rejected.

Using the P-value approach:

The p-value for two-tailed, "df=19" degrees of freedom, "t=-3.354102" is "p = 0.003333," and since "p=0.003333<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu" is different than 280, at the "\alpha = 0.05" significance level.